OpenStudy (anonymous):

if 30 students are asked to pick a number between 1 and 60, what is the probability at least 2 will choose the same number?

OpenStudy (anonymous):

I'm really rusty with probability ... here is my attempt :) Total Number of possibilities for 30 students to pick any number from 1 to 60: T=60^30 Number of possibilities for 30 students to pick a number from 1 to 60, so that all their numbers are unique: U=60!/30! And the probability we are looking for is: P=1-U/T = 1 - (60!/30!)/(60^30)

OpenStudy (anonymous):

AT LEAST TWO a set up for "all are different"

OpenStudy (amistre64):

its might be easier to find the probability of "at most 1 student "

OpenStudy (anonymous):

is fiddlearound not correct?

OpenStudy (amistre64):

that mighta been worded wrong; but im thinking its complememnt

OpenStudy (anonymous):

fiddlearound is right.

OpenStudy (amistre64):

since its saying at least 2 can pick the same # im thinking it has to do "with replacement" ...

OpenStudy (amistre64):

...could this pass for a binomial probability as well?

OpenStudy (anonymous):

you compute the probabilty that they are all different, then subtract from 1. the probability they are all different is most succinctly written as \[\frac{\frac{60!}{30!}}{60^{30}}\]

OpenStudy (anonymous):

i think you can approximate this with poisson, so i guess it relates to binomial, but i forget how. good luck computing this number!

OpenStudy (amistre64):

fixed number of trials, two outcomes, probability stays the same and one other factor that determines it if i have it right in me head

OpenStudy (amistre64):

n= 30, p=1/60, q = 59/60 perhaps?

OpenStudy (anonymous):

but we don't have 2 outcomes for each trial - or maybe I'm missing something ...

OpenStudy (amistre64):

my thought, as wrong as it might be, comes to .08891

OpenStudy (amistre64):

well, if the outcome is to get the same number; then the other option is to not get the same number ... but thats onlyh if i see it right

OpenStudy (anonymous):

you are making me think too hard. i just recognize this as birthday problem. you can also compute via \[(1-\frac{1}{60})(1-\frac{2}{60})...(1-\frac{29}{60})\] i think

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=1+-+%2860%21%2F30%21%29%2F%2860^30%29 wolfram gets the number 0.999....

OpenStudy (anonymous):

the answer in the back of the book says the probability is almost 1

OpenStudy (anonymous):

yes it should be very very close to one. it is almost certain that two will pick same number

OpenStudy (anonymous):

Is that the answer they expect : "almost 1" ? :)

OpenStudy (amistre64):

lol .... good, then I was wrong :)

OpenStudy (anonymous):

yeah thats exactly the answer in the back of the book!

OpenStudy (anonymous):

oh look at fillearound wolfram link. you see that it is .999858...

OpenStudy (anonymous):

is the outside of the book fuzzy too?

OpenStudy (anonymous):

hahahahah